### v2

# A Survey on Human-AI Teaming with Large Pre-Trained Models

Abstract.
In the rapidly evolving landscape of artificial intelligence (AI), the collaboration between human intelligence and AI systems, known as Human-AI (HAI) Teaming, has emerged as a cornerstone for advancing problem-solving and decision-making processes. The advent of Large Pre-trained Models (LPtM) has significantly transformed this landscape, offering unprecedented capabilities by leveraging vast amounts of data to understand and predict complex patterns. This paper surveys the pivotal integration of LPtMs with HAI, emphasizing how these models enhance collaborative intelligence beyond traditional approaches. It examines the potential of LPtMs in augmenting human capabilities, discussing this collaboration for AI model improvements, effective teaming, ethical considerations, and their broad applied implications in various sectors. Through this exploration, the study sheds light on the transformative impact of LPtM-enhanced HAI Teaming, providing insights for future research, policy development, and strategic implementations aimed at harnessing the full potential of this collaboration for research and societal benefit.
Human-AI Teaming, Artificial Intelligence, Large Pre-Trained Models, Large Language Models
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Computing methodologies Artificial intelligenceccs: Human-centered computing Human computer interaction (HCI)
1. Introduction
The journey towards blending human intelligence with advanced technology spans several centuries, beginning with inventive attempts to simulate human intellect in machines. One notable example from the 1770s is the "Mechanical Turk," a machine that appeared to play chess autonomously but was, in fact, operated by a person concealed inside it (Clark et al., 1999). This early endeavor, while not "Artificial Intelligence" (AI) in the modern sense, reflects the enduring fascination with creating technology that can mimic or complement human cognitive abilities. The formal realization of integrating human-like intelligence into machines culminated in 1956 at Dartmouth College, marking the official birth of Artificial Intelligence as a research field (Russell et al., 2021). As AI technology has advanced and become more prevalent over the last decade, researchers have identified limitations within purely automated systems (Mittelstadt et al., 2016; Goodfellow et al., 2015; Rodrigues, 2020) leading to renewed focus on augmenting AI with human expertise, aiming to harness the best of both worlds (Fig. 1(a)). This approach seeks to enhance AI applications and foster a symbiotic relationship between human and artificial intelligence.
Figure 1. A graphical depiction of the number of relevant articles submitted from 2013-2023. (a) illustrates the increasing engagement of research communities in collaborations between humans and AI over the years. (b) observes a notably sharp increase in submissions related to human-AI and large-scale models attributed to the recent emergence of large models.
Navigating the rapidly advancing AI landscape, the integration of human cognition with Large Pre-trained Models (LPtM), including Large Language Models (LLM) (Brown et al., 2020; Touvron et al., 2023) and Large Vision Models (LVM) (Kirillov et al., 2023; Radford et al., 2021; Jia et al., 2022), has initiated a transformative shift. These models are "pre-trained" on vast datasets before being fine-tuned for specific tasks. This paradigm has opened new avenues for collaborative problem-solving and decision-making. LPtMs are shaped and refined by human expertise in terms of ethical guidance, creativity, and contextual understanding. In turn, AI amplifies human capabilities by processing data at large scales, offering insights and augmenting decision-making. This paper undertakes an extensive survey of Human-AI teaming, analyzing the complex interactions between human agents and sophisticated pre-trained models across various fields. Our study aims to uncover the progress made, confront the challenges, and understand the implications of this evolving partnership.
The primary contribution of this survey is to provide an in-depth perspective on the complex and diverse landscape of Human-AI teaming with large pre-trained models. We explore this synergy by studying Human-AI collaboration that refines AI model behavior, effective HAI systems, safety concerns, and sector-specific applications. Our goal is to inform future research, influence policy-making, and guide development strategies, ultimately driving an effective and ethical integration of Human-AI collaboration across various facets of society.
Figure 2. A representation of the scope of this survey. (a) We cover four broad categories - Model improvements, Effective HAI systems, Safe and Trustworthy AI, and their Applications - in the human-AI domain, along with the recent developments in them with the help of LPtM. The overlap between the four petals illustrates the fact that the four categories are not mutually exclusive; the articles may belong to two or more categories at the same time. (b) The heatmap shows a general trend of the number of articles co-existing within subcategories, including the articles discussing large models within each subject area. The overlap between the applications is not analyzed, given their specific and varied nature. AI: Artificial Intelligence, HAI: Human-AI, UI/UX: User Interface/User Experience.
1.1. Scope of the Survey
Our survey focuses on the articles describing the developments of human-AI teaming through the years and how the introduction of large pre-trained models is reforming the field. The search for articles was conducted on Google Scholar using keywords such as "Human-AI", "Human-AI teaming", and other similar terms. For research focusing on the collaboration between humans and AI involving large models, additional keywords, including "large models", "large language models", and "large vision models" were utilized. Since human-AI teaming and large models have only recently gained a lot of interest from the research community given the recent popularity of large pre-trained models (Fig. 1(b)), some portion of studies cited in this survey could also be from arXiv preprints. We then sample the articles according to the structure desired for this survey, i.e., primarily covering Human-AI teaming and its large pre-trained model counterparts in better model training, effective human-AI joint systems, safe and secure HAI teaming, and their applications (Fig. 2(a)).
In order to show support for the theme of our survey article, we try to implement human-AI teaming with the help of large language models to prepare this review. For the human part, the authors survey the existing literature about human-AI teaming, collect a subset of articles according to the desired structure, organize the survey into relevant sections, study submission statistics and put together tables and visualizations, and finally, prepare the first complete draft of the manuscript. Using the AI part in human-AI teaming, we ask ChatGPT-4 (OpenAI, 2023) for its feedback regarding the language flow of each subsection, placement of the citations, and revising the language wherever necessary. This helps in the ease in identifying the missing elements and making necessary revisions. The final manuscript is, therefore, a product of human and AI collaboration.
1.2. Outline of the Survey
In each segment of our study, we commence by delineating traditional methodologies employed in human-AI collaboration, subsequently delving into the contributions of extensive pre-trained models in the fulfillment of these tasks. Section 2 explores the involvement of incorporating human expertise into the AI model training cycle, fostering a cooperative relationship between humans and AI in active learning scenarios, enhancing learning through human feedback, and involving human experts in the thorough evaluation of machine learning models (Hu and Sadigh, 2023; Settles, 2009; Ziegler et al., 2019). Section 3 delves into understanding the scope of effective Human-AI joint systems, optimizing user interfaces (UI) and system architectures emerges as a crucial area. It covers a range of topics under innovative UI designs and system structures aimed at boosting efficiency and user engagement in collaborative settings (Lemaignan et al., 2017; Fan et al., 2022).
A paramount concern in the realm of Human-AI collaboration is the safety, security, and trustworthiness of AI systems. Section 4 of our comprehensive analysis encompasses multiple dimensions: mitigating algorithmic biases to uphold fairness, assessing the impact of Human-AI collaboration on workers’ autonomy, well-being, and job satisfaction, examining economic repercussions on employment and wages, addressing data privacy and security concerns, building trust in AI systems, and navigating the legal and regulatory frameworks governing Human-AI interactions (Konstantis et al., 2023; Hemmer et al., 2023; Braun et al., 2021). These factors are integral to creating an ethical and responsible environment for AI’s integration into human-centric workflows. Furthermore, section 5 survey explores the broad spectrum of Human-AI teaming applications across various sectors. We investigate the unique challenges and opportunities in fields such as healthcare (Bienefeld et al., 2023), autonomous vehicles, surveillance systems (Hauptman et al., 2023), gaming (Frans, 2021), education, and accessibility. Understanding the specific characteristics and benefits of Human-AI collaboration in these areas is key to influencing its future direction and maximizing its societal impact.
2. AI Model Improvements with Human-AI Teaming
In this section, we delve into the process of training AI models, a critical stage that underpins their effectiveness and efficiency. Drawing upon the framework presented in Maadi et al. (Maadi et al., 2021), we explore the integral role of human-AI interactions in model training. This framework delineates three interdependent phases: 1) Data control, where data is pre-processed and produced under stringent quality standards; 2) Machine Learning (ML) modeling and execution, guided by human expertise; and 3) ML evaluation and refinement. Our discussion primarily centers on the second phase, highlighting its pivotal role in shaping models that fulfill human-centric goals of effectiveness and efficiency. This section covers four key approaches: Human in the Loop, where human judgment directly influences AI training; Active Learning, focusing on AI systems that efficiently incorporate human input; Reinforcement Learning from Human Feedback, where AI behavior is shaped by human feedback; and Human Evaluated Machine Learning, emphasizing the role of human evaluation in refining AI models. Together, these subsections provide insights into how human-AI collaboration can enhance the training process, leading to more effective and reliable AI models.
{tblr}
width = colspec = Q[m,250]Q[m,350]Q[700], row1 = c, column2 = m, column3 = t, cell21 = r=2c, cell41 = r=3c, cell71 = r=6c, cell131 = r=6c, vlines, hline1,19 = -0.08em, hline2,4,7,13 = -, hline3,5-6,8-12,14-18 = 2-3, Human-AI Topics & Subtopics Articles Cited
Section 2: AI model improvements with human-AI teaming Human in the loop (Whittlestone et al., 2019), (Hu and Sadigh, 2023), (Memmert and Tavanapour, 2023), (Freeman et al., 2014), (Settles, 2009), (Pries et al., 2023), (Lu et al., 2022), (Bao et al., 2018), (Lertvittayakumjorn et al., 2020), (Yao et al., 2023), (Christiano et al., 2017), (Ibarz et al., 2018), (Lee et al., 2023), (Kaelbling et al., 1996), (Sutton et al., 1999), (Ouyang et al., 2022), (Ziegler et al., 2019), (Nakano et al., 2021), (Rafailov et al., 2023a), (Zhang et al., 2021), (Johnson and Vera, 2019), (Mirchandani et al., 2023), (Zha et al., 2023), (McNeese et al., 2021)
Human evaluation in AI (Bansal et al., 2019), (Zhou and Chen, 2019), (AV et al., 2022), (Chhibber et al., 2022), (Hu and Sadigh, 2023), (Tabrez et al., 2020), (Correia et al., 2023), (Clark et al., 2021), (Ouyang et al., 2022), (Lu et al., 2023), (Chen et al., 2024), (Chiang and Lee, 2023), (Uchendu et al., 2023)
Section 3: Effective human-AI joint systems Improving user interfaces for effective teaming (Zhang et al., 2021), (Hauptman et al., 2023), (Bryan Wang, 2023), (Qingxiao et al., 2023), (Mingming et al., 2023), (Wu et al., 2022), (Kozlowski and Ilgen, 2006), (Salas et al., 2017), (Zhang et al., 2023), (Chhibber et al., 2022), (Clark et al., 2016), (Wang et al., 2019), (van den Bosch et al., 2019), (Flathmann et al., 2021), (Fan et al., 2022), (Wei et al., 2023), (Lemaignan et al., 2017)
Effective human-AI collaboration (Wei et al., 2023), (Wu et al., 2022), (Pflanzer et al., 2022), (Dubey et al., 2020), (Liu et al., 2023a), (Ngo et al., 2023), (Zhang et al., 2023), (Munyaka et al., 2023), (Gopinath et al., 2022), (Sharifi Noorian et al., 2022), (Rafailov et al., 2023b), (Nakano et al., 2021), (Ouyang et al., 2022), (Ziegler et al., 2019), (Zhou and Chen, 2019), (Henry et al., 2022), (McNeese et al., 2021), (Lv et al., 2021), (Wang, 2023), (Veselovsky et al., 2023), (Pal, 2023), (Memmert and Tavanapour, 2023), (Hu and Sadigh, 2023), (Ahmad et al., 2023), (Fuchs et al., 2023), (Bansal et al., 2021), (Chakraborti et al., 2017), (van den Bosch et al., 2019), (Liu et al., 2023b), (Neerincx et al., 2018), (Nikolaidis et al., 2017), (Michelle Zhao, 2022), (Kaptein et al., 2016), (Rastogi et al., 2023), (Sharma et al., 2023), (Chakrabarty et al., 2023)
Compatibility of human-AI systems (Weisz et al., 2021), (Chhibber et al., 2022), (Bau et al., 2019), (Wang, 2023), (Eloundou et al., 2023), (Harrer, 2023), (Niu et al., 2020), (McNeese et al., 2021), (Lyons et al., 2019), (Hou et al., 2023), (Bansal et al., 2021), (Wu et al., 2022), (Veselovsky et al., 2023), (Henry et al., 2022), (Liu et al., 2023b), (Munyaka et al., 2023), (Memmert and Tavanapour, 2023), (Chang et al., 2023)
Section 4: Safe, secure and trustworthy AI Algorithmic bias and fairness (Green and Chen, 2019), (Stahl et al., 2021), (Conlon et al., 2023), (Chhibber et al., 2022), (Morrison et al., 2023), (Kirk et al., 2021), (Huang et al., 2023), (Meyer et al., 2023), (Kotek et al., 2023), (Li et al., 2023), (Gallegos et al., 2023), (Ohi et al., 2024), (Bi et al., 2023)
Worker autonomy and well being (Chhibber et al., 2022), (Konstantis et al., 2023), (Pal, 2023), (Despotovic and Bogodistov, 2024), (Wang, 2023), (Woodruff et al., 2023)
Effect on wages and jobs (Hemmer et al., 2023), (Chowdhury et al., 2022), (Chhibber et al., 2022), (Pal, 2023), (Ashktorab et al., 2020), (Despotovic and Bogodistov, 2024), (Wang, 2023), (Walkowiak and MacDonald, 2023)
Data privacy and security (Ezer et al., 2019), (Yin et al., 2019), (Kaissis et al., 2020), (Samyuktha et al., 2022), (Lepri et al., 2021), (Hacker et al., 2023), (Ullah et al., 2023), (Gupta et al., 2023), (De Angelis et al., 2023), (Thapa and Adhikari, 2023), (Sebastian, 2023)
Trustworthy AI and accountability (Hou et al., 2023), (Pflanzer et al., 2022), (Caldwell et al., 2022), (Wang, 2023), (Kim et al., 2023), (Bansal et al., 2019)
Law and public policy (Jobin et al., 2019), (Cath et al., 2018), (Stahl et al., 2021), (Binz et al., 2023), (ThankGod Chinonso, 2023), (Walsh, 2023), (Marcos and Pullin, 2023)
Section 5: Applications Healthcare (Henry et al., 2022), (Bienefeld et al., 2023), (Memmert and Bittner, 2022), (McKinney et al., 2020), (Budd et al., 2021), (Choudhury and Asan, 2022), (Cai et al., 2019), (Lyu et al., 2023), (Kung et al., 2023), (Johnson et al., 2023)
Autonomous vehicles (Atakishiyev et al., 2021), (Lv et al., 2021), (Zhou and Chen, 2019), (Fuchs et al., 2023), (Yang et al., 2024), (Park et al., 2024), (Cui et al., 2024), (Wen et al., 2023), (Wang et al., 2023a), (Wang et al., 2023b)
Surveillance and security (Pazho et al., 2023), (Killcrece et al., 2003), (Hauptman et al., 2023), (Guo et al., 2018), (Chen et al., 2023), (Jain, 2023)
Games (Frans, 2021), (Xu et al., 2023), (Sobieszek and Price, 2022), (Akata et al., 2023), (Todd et al., 2023), (Värtinen et al., 2022), (Frans, 2021), (Hu and Sadigh, 2023)
Education (Nwana, 1990), (VanLehn, 2011), (Whitaker et al., 2013), (Kasinathan et al., 2017), (Eicher et al., 2018), (Hartle and Kaczorowski, 2019), (Dikli, 2006), (Silvia Milano, 2023), (Extance, 2023), (Wang and Demszky, 2023), (Hellas et al., 2023), (Chang et al., 2023)
Accessibility (Kumar and Jain, 2022), (Ozarkar et al., 2020), (Khan et al., 2020), (Wen et al., 2021), (Ghazal et al., 2021), (Gadiraju et al., 2023), (Taheri et al., 2023)
Table 1. A tabular representation of the four broader focus topics covered in this survey, with their relevant subtopics. Each category contains its cited articles for the ease of reader reference.
2.1. Human in the Loop
In the ‘Human in the Loop’ (HITL) approach, we explore the synergistic collaboration between human intellect and AI capabilities. This collaboration is especially critical in addressing AI’s limitations in self-learning and humans’ challenges in processing vast data within time constraints. The interplay of human and AI strengths leads to more robust and ethically sound AI systems. For instance, during the model design phase, human insights on ethical considerations, as highlighted by Whittlestone et al. (Whittlestone et al., 2019), enrich AI perspectives beyond mere technical aspects. Conversely, in the model training and execution phase, the directional flow of information between humans and AI plays a pivotal role. Strategies such as InstructRL, as proposed by Hu and Sadigh (Hu and Sadigh, 2023), demonstrate the effectiveness of NLP instructions from humans in guiding AI, enhancing the model’s explainability and alignment with our expectations. Similarly, a research on AI-assisted brainstorming illustrates how AI’s engagement with human ideas can foster greater creativity and innovation (Memmert and Tavanapour, 2023).
The diverse roles of humans, ranging from providing ethical oversight in the design phase to contributing to quality control and data labeling during model training and execution, are crucial for optimizing AI performance. These roles not only bridge gaps in AI knowledge but also ensure that AI systems are aligned with human values and needs.
As we proceed, we will delve deeper into two specific HITL machine learning approaches: Active Learning, where AI models dynamically integrate human input to refine learning, and Reinforcement Learning from Human Feedback (RLHF), which leverages human feedback to shape AI behavior. These sections will illuminate the intricate ways in which human involvement is woven into the fabric of AI development, showcasing the multifaceted nature of Human-AI collaboration.
2.1.1. Human-AI Collaboration with Active Learning
Unlike traditional machine learning paradigms, active learning involves an iterative process where a model selectively identifies data requiring labeling, optimizing system performance with minimal training data (Freeman et al., 2014). Particularly, this approach is crucial for fine-tuning large pre-trained models, especially transformer-based ones, which require substantial labeled data for specific user scenarios. Active learning efficiently harnesses human expertise to pinpoint areas of uncertainty, enabling more targeted training with less annotation effort (Settles, 2009). The active learning workflow begins with an unlabeled dataset and a pre-trained model. The model predicts labels for each sample, outputting confidence levels (Fig. 3). When predictions fall below a quality threshold, human annotators step in for manual annotation. This iterative process of re-training the model with new labeled data continues until satisfactory confidence levels are achieved. Human AI teaming plays a vital role in efficient data labeling and continuous model improvement.
Figure 3. A visual representation of an active learning workflow. If the model predictions do not pass the required confidence value thresholds required for a task, the data is sent to humans for manual annotations, further improving the model by retraining it with the now-labeled data.
Recent research in this area has introduced novel methods to leverage human expertise effectively. For instance, Pries et al. (Pries et al., 2023) propose a method for efficient data labeling using precise distance measurements to filter data samples, streamlining the presentation of data to experts. Lu et al. (Lu et al., 2022) discuss human-guided interventions for continuous model improvement, focusing on dissociating biases. Bao et al. (Bao et al., 2018) explore the transformation of human-annotated rationales into continuous attention mechanisms, enhancing the learning of domain-invariant representations. Additionally, Lertvittayakumjorn et al. (Lertvittayakumjorn et al., 2020) present a generalizable approach for debugging deep text classification models with human input, applicable to larger models. Finally, Yao et al. (Yao et al., 2023) propose an explainable-generation active learning framework for simpler tasks, potentially improving the sampling process in data augmentation.
2.1.2. Reinforcement Learning from Human Feedback
Reinforcement Learning from Human Feedback (RLHF) is a paradigm that integrates human insights into the training process (Christiano et al., 2017; Ibarz et al., 2018). This approach enhances the collaboration between humans and AI systems, especially during the data control phase — a critical stage for the success of large pre-trained models. In RLHF, the human role extends beyond mere data provision to actively shaping the AI’s learning trajectory. This human-AI symbiosis is particularly effective in mitigating biases in training data, leading to fairer AI outcomes. Additionally, the human touch in RLHF makes AI decisions more transparent and comprehensible, thereby fostering greater trust among users. The adaptability of RLHF models to user preferences highlights the significance of human understanding in complex, contextual situations, enabling AI systems to be more aligned with ethical standards and personalization needs.
However, the implementation of RLHF is not without its challenges. Ensuring the consistency and accuracy of human feedback, effectively scaling this approach for large datasets, and harmonizing human insights with algorithmic processes are areas that require careful consideration (Lee et al., 2023). Overcoming these challenges is essential to maximize the benefits of RLHF, paving the way for AI models that are not only technically proficient but also deeply attuned to human values and needs.
Reinforcement Learning Enhanced by Human Feedback
As an active research area of machine learning, Reinforcement Learning (RL) focuses on enhancing agent behaviors to make decisions by performing actions in an environment and receiving feedback as rewards or penalties. Due to the success of RLHF in large models (LMs) and the similar interaction in HAI teaming, researchers are now striving to figure out how to leverage RLHF for further improvement on HAI frameworks and even put forward more flexible ones.
Despite various theories and empirical models behind RL, as mentioned in (Kaelbling et al., 1996), RL frameworks can be abstracted as a continuous interaction between agents and their surrounding environment, in which they take actions individually or cooperatively given their states and state transitions to maximize long-term rewards. Traditionally, given a designated reward model, if the accumulated long-term reward is exactly or approximately converged under limited horizons, the agent’s behaviors and the corresponding dynamics (Sutton et al., 1999), like policy $\pi(a|s)$, Q function $Q(s,a)$, etc. are more likely to converge. Only after OpenAI ’s introduction of instructGPT (Ouyang et al., 2022) to illustrate how a reward model trained by high-quality data from labelers can optimize the LM’s ranking of generative content did research communities realize RL’s critical role in improving LM performance. Since labelers are asked to label this data in multiple turns to yield the most suitable outputs, this interaction pattern under RL formalization is considered RLHF. Furthermore, this working pattern can be widely applied in similar scenarios within feedback-acquiring processes. In the HAI teaming model, once human instructions are quantitatively converted to a reward model by ranking or scoring, the interaction between human and AI system agents can be formalized and optimized by RLHF.
Design of RLHF
In exploring the design of RLHF, a key challenge emerges: aligning human instructions with RL agent behaviors to ensure that these instructions effectively guide and evaluate agent actions. This alignment, complex in nature, has been a pivotal focus in recent large language model (LLM) research, contributing significantly to the advancement of LLMs and their applications. This section reviews the development of RLHF from its early stages in 2019 to its current evolution.
A notable early work in RLHF is by Ziegler et al. (Ziegler et al., 2019), where language models were trained to improve text generation and summarization based on human preferences. However, this approach, limited by its reliance on a single-run process and restricted data, faced challenges in generalization. In contrast, OpenAI’s instructGPT (Ouyang et al., 2022) introduced a three-phase training paradigm: initial training with supervised policy via demonstration data, training a reward model with labeler feedback, and optimizing the ranking policy using RL against the reward model. This methodology provides a robust baseline and ensures that GPT models evolve with continuous human feedback, leading to the creation of more sophisticated and responsive models, such as the web browser-assisted question-answering agent by Nakano et al. (Nakano et al., 2021).
Despite these advancements, RLHF remains a complex and often unstable process. To address this, Stanford’s recent research has introduced the Direct Preference Optimization (DPO) training model (Rafailov et al., 2023a). DPO simplifies policy training by directly mapping reward functions to optimal policies, circumventing the need for intricate reward maximization constraints. This approach not only surpasses traditional RLHF in controlling sentiment in generations but also enhances response quality in tasks like summarization and single-turn dialogue. Its simplicity in implementation and training offers a promising direction for further simplifying RLHF processes.
New Human-AI Teaming Paradigms in RLHF
The integration of Reinforcement Learning from Human Feedback (RLHF) into Human-AI (HAI) processes presents unique challenges due to the intricate steps and components involved, as evidenced by the work of Zhang et al. (Zhang et al., 2021), and Johnson and Vera (Johnson and Vera, 2019). Recognizing these complexities, recent research has shifted focus towards enhancing collaboration within the closely-coupled phases of HAI, with the aspiration of exploring further extensions in the future. A notable gap in current RLHF models is the communication barrier between AI and humans, highlighting the need for more advanced frameworks to bridge this divide.
A significant contribution to human-agent coordination is the instructRL model by Hu and Sadigh (Hu and Sadigh, 2023), which employs natural language instructions to align RL agents with human-preferred policies. This model leverages the embedded knowledge of pre-trained models, enabling AI agents to collaborate more effectively with humans across various tasks, including gaming, robotics, and control decision-making. InstructRL’s adaptability to different domains through knowledge transfer and meta-learning positions it as a versatile model for diverse HAI scenarios.
Research by Mirchandani et al. (Mirchandani et al., 2023) suggests that Large Language Models (LLMs), even without additional training, can function as proficient general sequence modelers due to their in-context learning abilities. This research explores the use of LLMs in robotics, particularly for extrapolating state sequences and prompting reward-conditioned trajectories, despite challenges like latency and compute costs. Chain-of-thought prompting further enhances LLM performance in complex reasoning tasks, as shown by Wei et al. (Wei et al., 2022). The Distillation and Retrieval of Online Corrections (DROC) system by Zha et al. (Zha et al., 2023) exemplifies the capability of LLMs to assimilate language feedback and distill knowledge from corrections, demonstrating effectiveness in novel settings.
However, McNeese et al.’s research (McNeese et al., 2021) provides a counterpoint to the notion of RLHF as a universal solution. Despite the success of NeoCITIES (McNeese et al., 2021) RL agents in simulated environments, the disparity between these environments and real-world complexities remains a significant hurdle. Furthermore, RL agents lack proficiency in non-verbal communication and still struggle with detailed verbal interactions, indicating a need for more sophisticated models that can fully comprehend both verbal and non-verbal human communication. These observations underscore the necessity for developing new HAI frameworks that go beyond existing RLHF models, calling for enhanced human integration to address the multifaceted nature of real-world environments.
2.2. Human Evaluation in AI
Human Evaluation in Artificial Intelligence (Human-Eval-AI) is an essential process that integrates human expertise with AI-driven insights. This collaborative approach leverages the unique strengths of human judgment to refine and guide AI decision-making. The objective is to develop models that are accurate and also resonate with the diverse and multifaceted nature of our world. Such collaboration is key to training robust models, enhancing decision-making processes, and fostering a symbiotic relationship to improve model training.
Figure 4. Evaluation training strategies based on Clark et al. (Clark et al., 2021). They employ training methods by giving example and comparison tests, along with their correct answers. This is followed by an explanation of why the answer is correct. This way, the evaluators gradually learn about how to distinguish between human and machine-generated texts.
2.2.1. Evaluation of Trust and Transparency
Establishing trust and transparency is crucial in Human-Eval-AI systems. Trust, inherently linked with transparency and understandability, aligns closely with how users perceive an AI system’s abilities and limitations. As highlighted by Bansal et al. (Bansal et al., 2019), it is essential for users to recognize the AI system’s potential for errors. This understanding enables users to develop a realistic mental model of the AI, crucial for anticipating possible failures or underperformance.
Zhou and Chen’s work (Zhou and Chen, 2019) further discusses this by recommending the integration of uncertainty measures and machine learning model performance into AI-informed human decision-making processes. They advocate for a trial-to-trial approach in decision-making, where users adjust their judgment boundaries based on their experience of uncertainty and AI interactions. Their proposed Uncertainty-Performance Interface (UPI) allows users to concurrently evaluate the quality and performance of machine learning models, fostering deeper trust in AI systems. Similarly, Av et al. (AV et al., 2022) emphasize the importance of incorporating intrinsic uncertainty into the collaborative process. They suggest a mechanism for human experts to offer feedback, either as corrections to current AI recommendations or insights on preferable aspects within the search space. This feedback is integrated into a Bayesian optimization model, aiming to align the model’s inferences with the expert’s domain-specific knowledge.
2.2.2. Approaches to Improve Human Evaluation
In Human-Eval-AI, human input remains essential, especially in instructing AI agents and refining AI decision-making processes. The study by Chhibber et al. (Chhibber et al., 2022) showcases the benefits of integrating human teachings into AI, not only enhancing AI performance but also building trust and influencing job delegation practices. In a similar vein, Hu and Sadigh’s instructRL (Hu and Sadigh, 2023) introduces a novel form of human-AI interaction through high-level natural language instructions, thereby improving AI interpretability and alignment with human intentions. Aaquib et al.’s research (Tabrez et al., 2020) on mental modeling in human-robot teaming reveals that mutual understanding between humans and AI significantly boosts team performance, advocating for AI systems capable of adapting to human mental models. Complementing this, Correia et al. (Correia et al., 2023) emphasize context in AI system design, proposing personalized user interactions for more relevant human-machine team actions and information.
Further exploring human evaluation, Clark et al. (Clark et al., 2021) examine how human evaluators can differentiate between texts produced by humans and models like GPT-2 and GPT-3. They suggest methods to enhance evaluation accuracy, including detailed instructions, annotated examples, and text comparisons (Fig. 4). Ouyang et al. (Ouyang et al., 2022) present a different approach, fine-tuning GPT-3 into ‘InstructGPT’ using human feedback for improved user instruction alignment. Additionally, Lu et al.’s Error Analysis Prompting (EAPrompt) (Lu et al., 2023) focuses on evaluating machine translations by LLMs, leveraging Chain-of-Thought reasoning to refine LLM evaluative capabilities. Chen et al’s work (Chen et al., 2024) shows domain specific evaluation of LLM in problem-solving. Chiang and Lee (Chiang and Lee, 2023) investigate LLMs for human-like text quality assessments in specific NLP tasks, highlighting similarities with human evaluations. Finally, Uchendu et al. (Uchendu et al., 2023) highlight the effectiveness of collaborative efforts in enhancing detection capabilities, offering a unique perspective on human collaboration in evaluations.
3. Effective Human-AI Joint Systems
This section concentrates on achieving efficiency in the joint human-AI (HAI) systems. The focus is on merging the strengths of different HAI systems to ensure optimal performance while maintaining compatibility between the two (Fig. 5).
3.1. Improving User Interfaces for Effective Teaming
The following module discusses enhancing human-AI collaboration through improved user interfaces and data processing. It highlights AI’s role in team adaptation, personalization, conversational dynamics, and User Interface/Experience (UI/UX) design. The discussion will be focused around AI’s influence on team collaboration and trust, its facilitation of learning and feedback via dialogue, and the significance of ethical, effective UI/UX design in supporting human collaboration.
Figure 5. The chart shows the topics discussed to understand effective human-AI joint systems in this survey. The topics broadly discuss better interfacing, collaboration, and compatibility of human-AI systems.
3.1.1. Adaptation and Personalization
The integration of AI agents into human teams has historically been limited by their narrow adaptability and personalization capabilities. Traditional AI systems struggled to understand human subtleties and adapt to team dynamics, being confined to specific tasks without the ability to learn or adjust, thereby reducing their effectiveness in collaborative settings (Zhang et al., 2021).
The introduction of large pre-trained models has significantly revolutionized AI’s adaptability and personalization in teamwork. These models, trained on extensive datasets and utilizing sophisticated algorithms, have a markedly improved capacity to comprehend and adapt to team changes. Research demonstrates that LLM-equipped AI agents can fine-tune their autonomy, aligning with human team dynamics and operational needs more seamlessly (Hauptman et al., 2023). LLMs have advanced conversational capabilities, enabling more natural, context-aware interactions and supporting more complex, multi-party engagements (Bryan Wang, 2023; Qingxiao et al., 2023). Their application in mobile UI and UX evaluation also showcases their role in personalizing interactions and enhancing team collaboration (Mingming et al., 2023). The practice of chaining LLM prompts underscores their potential to customize Human-AI collaboration, ensuring greater control and transparency in the process (Wu et al., 2022).
3.1.2. Conversational Interaction
In Human-AI Teams (HATs), effective communication is as crucial as in human teams, with the conversational capabilities of AI playing a pivotal role in maintaining trust, facilitating knowledge sharing, and ensuring smooth coordination. These capabilities are essential for both the cognitive processing of tasks and the emotional dynamics within the team, impacting decision-making and situational awareness (Kozlowski and Ilgen, 2006; Salas et al., 2017). Quick and appropriate responses from AI can enhance human members’ confidence in their AI partners, leading to better teamwork (Zhang et al., 2023).
The proactive sharing of information by AI, crucial for building trust and promoting collaboration, has seen significant strides with the development of Teachable Conversational Agents. These agents improve text classification tasks by learning from dialogue-based teaching and feedback, adjusting their focus based on interaction, which can range from acquiring new knowledge to refining existing information based on the conversation’s context (Chhibber et al., 2022). The perception of these agents’ likability and human-likeness is key to their effectiveness in communication (Clark et al., 2016). Voice-based interactions, supported by Natural Language Processing (NLP), offer substantial benefits for both verbal and written exchanges in HATs (Wang et al., 2019).
Research efforts have also been made to create a "Shared Vocabulary" (Taxonomy Model) for HATs, aiming for a unified set of terms and concepts accessible to both humans and AI. Furthermore, the "Communication and Explanation" framework (Communication Model) suggests equipping AI with the skills and training necessary for clear communication and explanations of their actions and reasoning, improving team understanding and efficiency (van den Bosch et al., 2019). This approach not only fosters mutual comprehension but significantly boosts team performance.
3.1.3. UI/UX
The evolution of User Interface (UI) and User Experience (UX) design in Human-AI Joint Systems is crucial in defining the quality of collaborative experiences. Historically, UI design principles have focused on ensuring user-friendly and ethical interactions between humans and AI systems. These ethical concerns, particularly about their impact on individuals and society, have become more pressing as AI’s role in human interactions expands (Flathmann et al., 2021). The complexity of such interactions demands rigorous usability testing, requiring a keen eye for detail and a profound understanding of human-computer interaction principles (Fan et al., 2022). Traditional methodologies have highlighted the importance of clear communication and the synchronization of AI in UX evaluations, allowing AI to articulate its findings in usability testing and to determine effective modes of collaboration, whether they occur in real-time or are delayed.
The recent introduction of large models has significantly transformed UI/UX design practices within Human-AI Joint Systems. Innovations such as the two-stage frameworks, with ChatIE being a prime example, have redefined zero-shot Information Extraction (IE) tasks into multi-turn question-answering sequences, marking a significant advance in the field. These models, including ChatIE, have outperformed numerous established models across various datasets (Wei et al., 2023). Further, studies involving human subjects in varied social settings have shown that these advanced approaches not only improve the quality of interactions but also lead to more engaging and emotionally gratifying user experiences. This reflects the burgeoning capability of artificial intelligence to adopt more human-like behaviors and seamlessly integrate into social contexts (Lemaignan et al., 2017). This progression highlights a significant shift in the landscape of UI/UX design in Human-AI systems, where the deployment of large-scale models plays an increasingly central role in enhancing user experience and interaction fidelity.
3.2. Effective Human-AI Collaboration
In this section, we categorize and discusses different interaction methods for human-AI collaboration. We divide these methods into 3 subsections: Human Helps AI, AI Helps Human, and their combination. Human Helps AI focuses on systems where the AI has the final say, and humans provide it with support and guidance along the way. Human oversight and intervention are also discussed in this subsection. AI Helps Human is the inverse; where a human produces the final output, using AI assistance along the way. The final part looks into systems where humans and AIs have more equal roles. Specifically, how the outputs of both humans and AI can be ‘combined’ to produce a higher quality output than either can produce alone.
3.2.1. Human Helps AI
The indispensable role of human input in the development and refinement of AI systems is well-documented, underscoring the symbiotic relationship between human expertise and AI capabilities. Human involvement is pivotal at various stages of AI system development, from initial data preparation to the fine-tuning of algorithms for enhanced precision and reliability.
Data annotation and preparation, as highlighted by Wei et al. (Wei et al., 2023), serve as the bedrock for training AI systems, ensuring they are equipped with accurate and contextually relevant data. The customization and refinement of AI algorithms, driven by human feedback and insights, further illustrate the central role of human expertise in AI development, enhancing the system’s accuracy and applicability (Wu et al., 2022). Ethical considerations and contextual understanding are also paramount, with Pflanzer et al. (Pflanzer et al., 2022) emphasizing the need for AI to operate within societal norms and comprehend complex human contexts—an aspect that purely algorithmic approaches may miss. The adaptation of AI systems to specific industry requirements showcases the importance of tailoring AI technologies to meet diverse application needs, thus enhancing their personalization and relevance (Dubey et al., 2020; Liu et al., 2023a; Ngo et al., 2023).
Interface design and communication strategies, as discussed by Zhang et al. (Zhang et al., 2023), play a crucial role in making AI tools both accessible and user-friendly, bridging the gap between sophisticated AI outputs and practical user interaction. The continuous evolution of AI systems necessitates ongoing human intervention and guidance to stay abreast of advancements in the field (Munyaka et al., 2023). In specialized domains like adaptive AI systems for driving, human behavioral data and feedback are utilized to simulate and incorporate diverse driving behaviors, demonstrating the integration of human subtleties into AI training for real-world applications (Gopinath et al., 2022). Hybrid human-AI teaming scenarios, where control is dynamically allocated between humans and AI based on situational awareness, further highlight the collaboration between human judgment and AI efficiency (Sharifi Noorian et al., 2022).
Recent advancements in large language models (LLMs) have also showcased the significant impact of human feedback on improving AI performance. Direct Preference Optimization (DPO), as introduced by Rafailov et al. (Rafailov et al., 2023b), represents a significant leap forward in the customization of language models. DPO focuses on aligning model outputs with human preferences through a direct optimization process, bypassing the complexities and limitations associated with traditional reinforcement learning approaches. Similarly, the work of Nakano et al. (Nakano et al., 2021) demonstrates how LLMs like GPT-3 can be fine-tuned using human feedback for tasks such as web-assisted question-answering. Furthermore, the instructRL framework introduces a method where AI agents, trained through reinforcement learning, are fine-tuned with human-provided natural language instructions, demonstrating the efficacy of intuitive human-AI collaboration (Hu and Sadigh, 2023). On the similar lines, Ouyang et al. (Ouyang et al., 2022) explore the benefits of fine-tuning LLMs with human feedback through the development of instructGPT models. The research of Ziegler et al. (Ziegler et al., 2019) complements these findings by applying reward learning, informed by human judgments, to tailor language models for specific tasks such as stylized text generation and summarization. This approach highlights the versatility of human feedback in enhancing AI models, enabling them to adapt to a wide range of stylistic preferences and content requirements.
3.2.2. AI Helps Human
This discussion delves into how AI systems, from traditional approaches to cutting-edge Large Language Models (LLMs), provide pivotal assistance across various domains, enhancing human capabilities and streamlining complex tasks. It traverses the evolution from conventional AI applications, such as medical diagnostic aids and vehicular control improvements, to the transformative potential of LLMs in data analysis, creative content generation, and decision-making, illustrating the dynamic progression of AI as an indispensable partner in human endeavors.
Initial discussions focus on traditional AI’s role in simplifying complex tasks, like transforming zero-shot information extraction into manageable question-answering formats, aiding humans in navigating large datasets effectively (Wei et al., 2023). Additionally, AI’s contribution to enhancing human decision-making, particularly under uncertainty, is explored through its applications in autonomous vehicles, emphasizing the need for transparency and trust in AI-human collaborations (Zhou and Chen, 2019). The use of ML-based systems in healthcare, such as TREWS for sepsis detection, exemplifies AI’s capacity to act as a "second pair of eyes," assisting clinicians with the voluminous clinical data for timely interventions (Henry et al., 2022). Experimental studies on team dynamics, incorporating both humans and AI agents, reveal the efficiency gains from AI integration, where AI handles tasks suited to its capabilities, thus optimizing team performance (McNeese et al., 2021). Furthermore, innovations in human-machine collaboration, such as intelligent haptic interfaces for vehicular control, showcase improvements in safety and efficiency during the transition from automated to manual driving (Lv et al., 2021).
More recently, impressive capabilities of LLMs in assisting humans, like chaining multiple LLM prompts offer a novel approach to complex task management, enhancing the human-AI collaborative experience through a transparent and controllable interface (Wu et al., 2022). LLMs’ role in augmenting, rather than replacing, human tasks is further detailed in Wang’s work (Wang, 2023), highlighting their utility in providing insights and suggestions for tasks that would otherwise be daunting for humans. Further, Veselovsky et al. (Veselovsky et al., 2023) touch on the balance required in leveraging LLMs to maintain the authenticity of human responses, underscoring the potential for high-quality but homogeneous outputs versus lower-quality but genuine human responses. Interdisciplinary collaboration is advocated as a means to shape the future of LLMs and AI-human interaction, promoting a proactive and global perspective (Pal, 2023). The impact of AI on human brainstorming processes, where systems like GPT-3 stimulate broader idea generation, is also examined, indicating AI’s potential to enhance cognitive processes and creativity in individual and collaborative settings (Memmert and Tavanapour, 2023). The application of ChatGPT in software architecture, aiding in the translation of architectural narratives into technical specifications, exemplifies AI’s role in enhancing productivity and fostering creative design processes (Ahmad et al., 2023). Furthermore, the examination of AI’s role in driving safety and autonomous systems, via a delegation manager that intelligently assigns control based on the situation and capabilities, illustrates the potential for AI to significantly improve operational safety and efficiency in hybrid human-AI systems (Fuchs et al., 2023).
3.2.3. Combining Human and AI Strengths
The essence of merging the strengths of humans and AI lies in leveraging the unique capabilities of each. It involves a careful balance where each compensates for the other’s limitations, aiming for a collaboration where the overlap in errors is minimized, thus maximizing the opportunity for mutual correction and improvement. Research indicates that while explanations of AI predictions can foster blind trust, they may not always promote the appropriate reliance needed in critical domains where human oversight is crucial (Bansal et al., 2021). Optimal collaboration requires an intricate understanding of each other’s capabilities, where humans excel in cognitive skills like context recognition, anticipation of team behaviors, and corresponding action support (Chakraborti et al., 2017).
However, recognizing and integrating these diverse forms of intelligence within Human-AI Teams (HATs) poses new challenges, necessitating an in-depth understanding of AI’s capabilities and limitations by human teammates (van den Bosch et al., 2019; Liu et al., 2023b). Initially, mismatches in mental models of AI’s abilities can lead to errors, but over time, as humans develop a more accurate understanding, they learn to effectively complement the AI’s functions. Advanced hybrid AI approaches, including neuro-symbolic frameworks, have been introduced to model human psychological states, aiding in the prediction and adaptation to human behaviors for enhanced cooperation (van den Bosch et al., 2019; Neerincx et al., 2018; Nikolaidis et al., 2017).
Adaptation in AI behavior, aimed at enhancing collaboration, involves learning from relevant human features and modifying behavior to align with human goals, thereby fostering a more intuitive and effective partnership (Michelle Zhao, 2022). Incorporating computational models of emotion into mental models allows AI agents to consider affective states in decision-making, enriching the collaboration with sophisticated, goal-oriented actions (Kaptein et al., 2016). Establishing a common vocabulary and shared understanding of goals, capabilities, and emotional states further solidifies the foundation of effective Human-AI collaboration (van den Bosch et al., 2019).
Recent research initiatives showcase practical applications of this synergy. AdaTest++, an augmented tool for auditing large language models (LLMs), exemplifies the collaborative effort between humans and AI in enhancing model reliability and understanding, facilitating the identification of model failures through a combined effort (Rastogi et al., 2023). In the realm of mental health support, the HAILEY system demonstrates how AI can augment human empathy with analytical insights, providing feedback-driven support for more empathetic and effective peer-to-peer conversations (Sharma et al., 2023). Additionally, the innovative integration of LLMs and diffusion models for creating visual metaphors highlights the creative potential unleashed by human-AI collaboration, where human-proposed linguistic metaphors are visually realized through AI-generated imagery, bridging the gap between conceptual ideas and visual representation (Chakrabarty et al., 2023).
3.3. Compatibility of Human-AI Systems
In human-AI joint systems, compatibility is crucial for ensuring that AI systems can seamlessly work alongside existing human technologies, thereby enhancing the overall efficacy of the system. As systems grow in complexity, system designers and engineers face the intricate challenge of ensuring harmonious interaction between humans and AI. This integration encompasses aspects such as communication protocols, security measures, scalability considerations, and workflow integration. When AI systems are tailored to fit within the existing workflows of human users, they facilitate smooth integration, avoiding disruptions or compromises to the system’s current functioning. In the context of human-AI systems, compatibility transcends technical alignment. It encompasses the adaptability of AI systems to existing human processes, the user-friendliness and learning curve associated with new AI tools, and the effectiveness of these compatible joint systems in achieving their intended results. This section delves into these key aspects, highlighting the importance of compatibility in the effectiveness and safety of human-AI systems.
3.3.1. Backwards Compatibility
In the realm of AI technologies, backward compatibility plays a pivotal role in ensuring the smooth integration of new or updated systems with pre-existing frameworks. This concept is crucial when embedding AI into legacy systems, as it involves careful consideration of existing tools, processes, and practices. Weisz et al. (Weisz et al., 2021) delve into how generative models contribute to the modernization of applications, emphasizing the necessity for AI systems to function cohesively within older infrastructures. In the context of traditional crowd work, Chhibber et al. (Chhibber et al., 2022) illustrate the significance of aligning AI tools with established workflows, showcasing the potential of generative models, like Generative Adversarial Networks (GANs) and autoencoders, to augment and refine datasets in legacy systems. Additionally, Bau et al. (Bau et al., 2019) present a novel approach using GANs to adapt image priors to specific image statistics in individual photographs, thereby tackling challenges in reproducing and manipulating high-level image attributes, as evidenced in various semantic editing tasks.
Discussing the broader compatibility challenges outside just the technological perspective, Wang (Wang, 2023) talks about the transformative effects of Large Language Models (LLMs) on the job market, underscoring the necessity of integrating these advanced technologies into existing economic structures without causing major disruptions. Eloundou et al. (Eloundou et al., 2023) examine the impact of LLMs on the US labor market, revealing that approximately 80% of workers could experience a significant change in their tasks, with 19% potentially seeing over half of their tasks affected. This ‘exposure’ refers to a 50% increase in task efficiency without compromising quality. Furthermore, Harrer et al. (Harrer, 2023) highlight the potential of LLMs in generative AI applications, emphasizing the need for responsible design, human oversight, and ethical considerations to prevent the spread of misinformation. They offer insights into how these models, when developed responsibly, can serve as efficient and trustworthy tools in sectors like healthcare, addressing technology, risks, and ethical issues.
3.3.2. Learning Curve and Effective Usability
In Human-AI Teaming (HAT), a critical factor is the ease and efficiency of interaction between humans and AI which significantly influences the learning curve and overall usability. McNeese et al.’s (McNeese et al., 2021) study talks about the team composition in HAT, revealing that AI’s ability to take a leading role in teams can enhance performance, though it often presents an initial learning curve for human participants. Lyons et al. (Lyons et al., 2019) discuss the importance of intuitive interfaces in fostering a partnership, rather than a tool-based relationship, between machines and humans. A pivotal aspect in this regard is the AI’s ability to explain its decisions, especially when those decisions may seem counter-intuitive to human collaborators. Hou et al.’s (Hou et al., 2023) research highlights that trust in AI is bolstered when humans understand the rationale behind AI decisions, fostering a sense of vulnerability and reliance on AI actions. This concept is further supported by Bansal et al. (Bansal et al., 2021), who find that AI explanations significantly increase human trust in the algorithm.
Exploring the potential of Large Models, Wu et al. (Wu et al., 2022) also explore that breaking down the complex human tasks into multiple LLM prompts and chaining them together helps with increased human-AI interaction towards improvement. The users improvised over the new ways of interacting with the LLMs by leveraging the sub-tasks to set calibrate model expectations and making new-unit tests for each submodule of the LLM chain. Veselovsky et al.’s (Veselovsky et al., 2023) study, however, suggests that even though prevalent usage of LLMs generating high-quality but homogeneous responses might not truly capture the real essence of human behavior, they can prove to be useful too. They analyze that human workers not using LLMs in their tasks produced lower quality (less essential) information than when allowed the use of LLMs. They, hence, given the rapidly evolving large model landscape, emphasize a need to constantly monitor the co-evolution of LLM and human teaming to get to some constructive contributions.
3.3.3. Evaluating the Effectiveness of Joint Human-AI Systems
Evaluating the efficacy of human-AI systems extends beyond technical performance, encompassing a holistic approach that includes technical, social, and psychological dimensions of human-AI collaboration. The study by Henry et al. (Henry et al., 2022) emphasizes the importance of user trust and system usability, suggesting that evaluation frameworks should assess how systems support user autonomy and are endorsed by domain experts to facilitate adoption. Similarly, Liu et al. (Liu et al., 2023b) highlight the need to understand perceptual differences between humans and AI, advocating for evaluations that measure how these systems leverage the complementary strengths of human and machine intelligence to improve joint decision-making processes. Munyaka et al. (Munyaka et al., 2023) delve into the impact of decision-making styles and AI identity disclosure on team dynamics and efficacy, underlining the significance of evaluating social dynamics and collaboration patterns within human-AI teams. Furthermore, Memmert et al. (Memmert and Tavanapour, 2023) explore the integration of AI in creative processes, pointing to the necessity of assessing both the cognitive contributions of AI and its effects on group dynamics, such as the potential for free-riding.
In terms of large models, Chang et al. (Chang et al., 2023) present a comprehensive theoretical framework for evaluating LLMs. This framework spans a wide array of metrics and scenarios, such as natural language processing, reasoning capabilities, and robustness. This study highlights the necessity of multifaceted evaluation frameworks that extend beyond task performance to include considerations like ethical implications and bias, which are crucial in human-AI systems. On the other hand, Veselovsky et al.(Veselovsky et al., 2023) adopt a more empirical approach by examining the practical use of LLMs in crowd work. Their evaluation method involves observing and adapting the behavior of crowd workers as they interact with LLMs, offering a practical perspective on Human-AI interaction. These studies collectively suggest that a comprehensive evaluation of human-AI systems should not only focus on accuracy or efficiency but also consider factors like trust, usability, collaboration dynamics, and the psychological impact on users, to ensure the development of systems that are effectively integrated into human workflows and decision-making processes.
4. Safe, Secure and Trustworthy AI
The integration of AI in professional domains has been both revolutionary and disruptive, significantly affecting employment and wage dynamics. In this section, we embark on a detailed examination of the multifaceted implications of these developments. First, we address the critical issue of algorithmic bias and fairness. By drawing on existing literature, we explore the complexities surrounding this challenge and investigate potential solutions. Next, we assess how AI impacts worker autonomy and job satisfaction, shedding light on both positive and negative effects. The concept of trustworthy AI is then analyzed, with a focus on the crucial aspect of accountability in human-AI interactions. Subsequently, we explore the intricate relationship between AI, law, and public policy, underscoring the challenges in developing regulatory frameworks that effectively balance ethical considerations with rapid technological advances. Our objective throughout this section is to provide an informed perspective on human-AI collaboration and its socio-economic implications, contrasting these new models with traditional paradigms.
Figure 6. Large Language Models often exhibit gender bias due to primarily being trained on large datasets of human language that reflect existing societal biases. Kotek et al. (Kotek et al., 2023) test this by probing LLMs to answer permutations of complex questions, as shown above. The answers and their explanation help the authors to quantify the bias.
4.1. Algorithmic Bias and Fairness
Algorithmic fairness is a cornerstone of contemporary AI, demanding the development and deployment of unbiased and non-discriminatory algorithms. This pursuit, though challenging due to the complexities in defining fairness, mitigating data biases, and ensuring algorithm transparency, has become a focal point for researchers and organizations (Green and Chen, 2019). Fairness in algorithms is inherently complex. For instance, biases in the training data can lead to skewed outputs, and inherent biases in algorithmic design can persist even with unbiased data. These issues are further complicated by the often opaque nature of algorithms and the lack of established accountability measures (Stahl et al., 2021). Several strategies have been proposed to address these challenges. Key among them are ensuring data representativeness and employing algorithms known for their fairness. Pre-deployment testing for biases, enhancing transparency, and establishing robust accountability mechanisms are equally crucial (Conlon et al., 2023). Chhibber et al’s (Chhibber et al., 2022) research sheds light on the role of user interaction in shaping algorithmic outcomes. It demonstrates how data quality, user engagement, and system localization impact fairness in machine learning systems. Additionally, Morrison et al.’s (Morrison et al., 2023) study discusses the value of causal explanations in AI, arguing that they foster trust, facilitate error correction, and support informed decision-making, thereby contributing to the enhancement of algorithmic fairness.
The incorporation of Large Language Models (LLMs) into various applications has intensified the need to address algorithmic bias and fairness. Kirk et al. (Kirk et al., 2021) have shown that biases in large language models, particularly those available ‘out-of-the-box’ like GPT-2, can lead to stereotypical and less diverse occupational associations for women, especially when intersecting with other categories like religion or ethnicity. Meanwhile, Huang et al. (Huang et al., 2023) reveal that a significant portion of code functions generated by LLMs can contain biases related to sensitive attributes such as age and gender, raising concerns about the fairness and integrity of software applications reliant on these models. Kotek et al. (Kotek et al., 2023) further investigate gender stereotypes within LLMs (Fig. 6), demonstrating that these models are significantly more likely to align occupations with stereotypical gender roles, sometimes amplifying biases beyond societal norms or official statistics. They find out that, on average, these LLMs are 6.8 times more likely to choose a stereotypical female answer when a female pronoun is present and 3.4 times for male answers and pronouns. Further, Meyer et al. (Meyer et al., 2023) explore the paradigm shift brought by conversational LLMs, discussing their potential to enhance academic work while cautioning against their inherent biases and accuracy issues.
Recent studies are also working on diverse strategies for mitigating bias in LLMs, emphasizing the importance of fairness across their applications. Li et al. (Li et al., 2023) review fairness strategies for both medium and large-sized LLMs, highlighting the need for tailored debiasing methods based on model size and training paradigms. Gallegos et al. (Gallegos et al., 2023) offer a detailed survey on bias evaluation and mitigation, categorizing techniques across model development stages and introducing taxonomies for bias metrics and mitigation methods. Ohi et al. (Ohi et al., 2024) propose a novel approach to counteract likelihood bias in LLM evaluations using few-shot examples for in-context learning, showing marked improvements in fairness. Bi et al. (Bi et al., 2023) focus on group fairness in social media, presenting a hierarchical schema for bias evaluation and a chain-of-thought mitigation method, GF-Think, to reduce bias. These contributions underline possible approaches to bias mitigation, aiming for ethical and equitable use of LLMs.
4.2. Worker Autonomy and Well Being
In the realm of AI-enhanced platforms, worker autonomy, well-being, and job satisfaction are pivotal, directly influencing engagement and productivity. Current AI practices sometimes neglect these factors, leading to dissatisfaction and turnover. Addressing this, several studies propose frameworks to improve these aspects in AI-centric work environments.
Chhibber et al. (Chhibber et al., 2022) demonstrate how teachable conversational agents in crowd-based applications can empower workers, enhancing control, personalized learning, and ownership, thereby boosting job satisfaction. Konstantis et al. (Konstantis et al., 2023) emphasize transparency and fairness in crowdsourcing platforms, advocating for clear decision-making and respectful treatment to improve worker experience. Furthermore, Pal (Pal, 2023) highlights the design of LLMs focused on worker well-being and autonomy, avoiding stress-inducing features and promoting engagement and learning.
Despotovic and Bogodistov (Despotovic and Bogodistov, 2024) highlight a trend in job seekers preferring roles that incorporate advanced AI, like ChatGPT, aligning with their identities and technological interests. This shift towards AI-interactive jobs indicates a broader preference for technologically engaging roles, impacting job satisfaction and autonomy. Complementarily, Wang (Wang, 2023) explores the paradox of LLMs in the job market, noting their role in both creating new opportunities and obsoleting certain jobs. This dual effect is key in assessing worker autonomy and well-being, as LLM-driven automation fosters more creative roles and autonomy but also raises concerns about job security and satisfaction. Integrating insights from recent research, it’s evident that while generative AI has the potential to automate menial tasks and foster more creative roles, it also poses challenges to worker autonomy, necessitating a balanced approach to AI deployment in the workplace (Woodruff et al., 2023).
4.3. Effect on Wages and Jobs
The influence of AI on labor, particularly wages and job security, is a subject of global debate. The discourse ranges from AI augmenting human capabilities and creating new jobs to concerns about task automation and worker displacement.
Hemmer et al. (Hemmer et al., 2023) suggest AI could enhance productivity by up to 40%, potentially leading to higher wages. In contrast, Chowdhury et al. (Chowdhury et al., 2022) point to AI’s role in both creating and displacing jobs, with an estimated 1.7 million new jobs and significant displacements in the U.S. by 2029. This highlights the necessity for proactive support for those affected by AI-driven changes. Chhibber et al. (Chhibber et al., 2022) discuss AI’s impact on the gig economy, noting potential job losses in areas like customer service, balanced by new opportunities in AI-related fields. Pal (Pal, 2023) foresees a growing LLM industry, suggesting increased job creation in AI R&D. Conversely, Wang (Wang, 2023) anticipates the displacement of up to 800 million jobs globally by 2030, advocating for protective legislation and equitable AI benefit distribution. Ashktorab et al. (Ashktorab et al., 2020) examine Human-AI collaboration in the gaming industry, where AI’s role in automating tasks like game testing could lead to job enrichment in creative roles.
The integration of Large Language Models has diverse implications for the job market. Despotovic and Bogodistov’s (Despotovic and Bogodistov, 2024) research reveals that job seekers show a preference for roles that resonate with their personal identity and involve frequent interaction with technologies like ChatGPT. This sees a shift in job market dynamics and the necessity for businesses to adapt their recruitment strategies. Wang (Wang, 2023) discusses the paradoxical effect of LLMs, highlighting both job creation and displacement across various sectors. This includes an examination of LLMs’ influence on routine tasks, specialized professions, and broader economic and societal aspects. Walkowiak et al. (Walkowiak and MacDonald, 2023) focus on the Australian workforce, quantifying the exposure to GenAI risks and potential job transformations, addressing privacy, cybersecurity, and ethical issues.
4.4. Data Privacy and Security
In the realm of human-AI collaboration, trust, particularly in data privacy and security, is foundational. Ezer et al. (Ezer et al., 2019) emphasize the critical role of trust engineering, advocating for algorithmic transparency and dynamic management to establish mutual trust in complex, information-imperfect environments. Yin et al. (Yin et al., 2019) further this by proposing a privacy-focused, human-centric approach, leveraging logistic regression and differential privacy to strike a balance between privacy and utility, achieving notable classification accuracy. In medical imaging, Kaissis et al. (Kaissis et al., 2020) highlight the imperative for secure, privacy-preserving AI through federated learning, addressing the privacy-utilization dichotomy to advance patient care. The synergy of AI and cybersecurity, as explored by Admin et al. (Samyuktha et al., 2022), showcases AI’s potential to fortify cyber defenses, underscoring the role of machine and deep learning in countering complex cyber threats and safeguarding data integrity. Lepri et al. (Lepri et al., 2021) advocate for a human-centric AI ethos, stressing the importance of privacy, accountability, and ethical decision-making in building secure, trustworthy AI systems.
Regulatory and privacy challenges for Large Generative AI Models (LGAIMs), including ChatGPT and GPT-4, demand targeted legal and technical solutions for safe use. Hacker et al. (Hacker et al., 2023) recommend a regulatory framework for LGAIMs focusing on transparency, risk management, and equitable treatment, particularly for high-risk uses. Ullah et al. (Ullah et al., 2023) propose PrivChatGPT, a model incorporating differential privacy for LLMs, highlighting the balance between privacy and utility. Gupta et al. (Gupta et al., 2023) discuss cybersecurity risks with GenAI, emphasizing the need for ethical practices and secure defenses against potential misuse. De Angelis et al. (De Angelis et al., 2023) identify LLMs as vectors for misinformation, advocating for policies to combat the "AI-driven infodemic" in public health. Thapa and Adhikari (Thapa and Adhikari, 2023) balance LLMs’ benefits against misinformation risks in biomedical research, calling for strict validation. Sebastian (Sebastian, 2023) focuses on privacy in AI chatbots, suggesting federated learning and data minimization to protect user information, reflecting increasing privacy concerns among LLM users.
4.5. Trustworthy AI and Accountability
Initially, conventional methods in trustworthy AI emphasized the importance of ethical principles, human oversight, and transparency. Hou et al. (Hou et al., 2023) discuss these principles, stressing the need for secure systems that identify risks while remaining under human control. Pflanzer et al. (Pflanzer et al., 2022) introduce the Agent-Deed-Consequence model as a traditional ethical algorithm framework, emphasizing human oversight in AI ethics. Caldwell et al. (Caldwell et al., 2022) highlight factors influencing trust in AI in cooperative environments, such as multiplayer gaming, underscoring the importance of environmental factors and peer influence in conventional trust-building.
The advent of LLMs introduces new dimensions to these conventional methods. In healthcare, a case study using LLMs (Wang, 2023) showcases the balance between patient engagement and AI over-reliance risks, underscoring the advancements in data sensitivity and patient communication. Kim et al. (Kim et al., 2023) demonstrate how LLMs contribute to managing uncertainty in human-AI collaboration, proposing a trial-to-trial approach to effectively communicate and estimate uncertainty. Furthermore, LLMs influence the formation of mental models, as discussed by Bansal et al. (Bansal et al., 2019). They highlight LLMs’ role in enhancing predictability and understanding of AI’s error boundaries, crucial for developing accurate mental models for better human-AI collaboration.
4.6. Law and Public Policy
Traditionally, AI legal and policy frameworks have focused on ensuring transparency, accountability, fairness, and bias mitigation. Pioneering work by Jobin et al. (Jobin et al., 2019), which analyzed AI guidelines globally, highlights the universal emphasis on ethical standards, transparency, and accountability. Similarly, research by Cath et al. (Cath et al., 2018) highlights the critical need for transparent and accountable AI, especially in decision-making processes. Stahl et al.(Stahl et al., 2021) propose six guidelines for AI regulation: Transparency, Accountability, Bias Mitigation, Human Oversight, Data Protection, and Public Education.
The integration of Large Language Models (LLMs) into the AI landscape underscores the need for updated legal frameworks to address new challenges in data use and AI-generated content regulation. Bender et al. (Binz et al., 2023) examine the ethical implications of LLMs, focusing on data usage and biases. An accredited law partner (Walsh, 2023) specializing in technology and intellectual property discusses U.S. court cases like GitHub Copilot’s use of open-source code, highlighting the legal intricacies of data attribution and compliance. This situation prompts a reevaluation of the datasets used for AI training and the adaptation of laws, particularly concerning copyrights and patents. Ekenobi et al. (ThankGod Chinonso, 2023) explore the legal impacts of deploying LLMs, with a focus on data protection and privacy. They commend ChatGPT’s privacy measures as a step towards mitigating potential risks. Additionally, Marcos and Pullin’s article (Marcos and Pullin, 2023) talks about the EU’s GDPR challenges, emphasizing the balance between technological growth and privacy rights. This necessitates robust regulatory frameworks that ensure LLM compliance with strict data protection standards, highlighting the importance of new laws in safeguarding privacy while fostering innovation.
5. Applications
This section highlights the role of Human-AI (HAI) Teaming and Large Pre-trained Models (LPtM) in revolutionizing fields such as healthcare, autonomous vehicle domain, surveillance, gaming, education, and accessibility. It highlights the symbiosis between human expertise and AI’s analytical power, emphasizing the importance of ethical AI development and user-centric design in maximizing benefits and inclusivity within society.
5.1. Healthcare
In the healthcare sector, recent researchers highlight the studies performed in enhancing patient diagnosis accuracy and fostering trust, emphasizing AI’s transformative influence in human-AI collaboration (Henry et al., 2022; Bienefeld et al., 2023; Memmert and Bittner, 2022). Advances such as the application of AI in breast cancer detection have markedly boosted diagnostic efficacy, sometimes even outperforming human experts, showcasing the remarkable potential of AI in healthcare (McKinney et al., 2020). These developments necessitate reevaluating AI’s role in patient care, addressing ethical considerations and practical challenges (Budd et al., 2021). Research focusing on healthcare practitioners’ views within the US highlights the critical factors of training, trustworthiness, and perceived AI risks, which are pivotal in its acceptance and application in clinical environments (Choudhury and Asan, 2022). Additionally, improvements in AI-powered medical image retrieval systems have been shown to augment both diagnostic capabilities and user trust, advocating for integrated systems that blend human insight with algorithmic accuracy (Cai et al., 2019). This partnership between AI technologies and healthcare professionals is vital for enhancing patient care outcomes and alleviating professional strain, albeit confronted by hurdles related to training and the accountability of AI systems.
Integrating Large Pre-trained Models such as ChatGPT into healthcare has shown promise in enhancing medical education, patient communication, and clinical decision-making. Lyu et al. (Lyu et al., 2023) demonstrated ChatGPT’s ability to simplify radiology reports for patients and healthcare providers, highlighting the potential for improved clinical education despite noting areas for accuracy and specificity enhancements. Kung et al. (Kung et al., 2023) found that ChatGPT scored at or near the passing threshold on the USMLE, suggesting its capability to support medical education and hinting at future roles in clinical settings. Johnson et al. (Johnson et al., 2023) further validated ChatGPT’s usefulness by showing its generally accurate responses to a range of medical questions, though recognizing the need for further refinement.
Figure 7. An example of how multimodal foundation models and large language models can be used for enhancing generalization and robustness in autonomous driving, as studied by Wang et al. (Wang et al., 2023b). They develop pixel/patch-aligned feature descriptors and latent space simulation, enriched with language modality, suggesting potential for optimizing the training and debugging processes for end-to-end learning-based control. The RGB images in this figure are taken from the KITTI dataset (Geiger et al., 2012) for representation purposes.
5.2. Autonomous Vehicles
Advancements in autonomous driving promise safer, AI-enhanced transportation, with significant milestones achieved in vehicle sensing and decision-making (Atakishiyev et al., 2021). Key to these advancements is understanding Human-AI collaboration, particularly during critical transitions between automated and manual driving modes, as explored by Lv et al. (Lv et al., 2021). This research highlights the importance of seamless transitions for safety, showing improvements in control and stability with a new hybrid driving model. Additionally, establishing trust in AI through transparent decision-making processes under uncertainty is crucial, as Jianlong and Chen (Zhou and Chen, 2019) discuss with a novel framework aimed at enhancing Human-AI collaboration. Fuchs et al. (Fuchs et al., 2023) further examine the managerial aspects of task delegation between humans and AI, proposing reinforcement learning to optimize collaboration and improve autonomous system performance, including driving.
Building on foundational studies, recent explorations into large language models (LLMs) signal a transformative era in autonomous driving technologies. Yang et al.(Yang et al., 2024) underscore the importance of LLMs for interpreting user commands in complex scenarios, emphasizing the necessity of model quality and prompt design. Park et al. (Park et al., 2024) introduce VLAAD, a multi-modal LLM that demonstrates proficient interpretive capabilities across varied driving situations, showcasing the potential of natural language models in enhancing decision-making processes. Cui et al. (Cui et al., 2024) propose a framework that integrates LLMs for improved decision-making in autonomous vehicles, aiming for personalized assistance and efficient operation. Wang et al. (Wang et al., 2023a) discuss using LLMs as behavior planners to ensure safety in decision-making by interpreting driving scenes and adjusting trajectory plans based on safety constraints. More generally, Wen et al. (Wen et al., 2023) highlight the role of Visual Language Models (VLMs) like GPT-4V in scene understanding and decision-making, noting superior performance but also pointing out existing challenges. Additionally, Wang et al. (Wang et al., 2023b) describe an end-to-end autonomous driving system using multimodal foundation models to enhance adaptability and robustness in diverse environments, leveraging pixel-aligned features and language-based latent space simulation for improved data augmentation and policy debugging (Fig. 7). Together, these studies highlight the transformative potential of large foundation models in advancing autonomous vehicle technologies, from improving interpretive capabilities to enhancing decision-making processes and data generation.
5.3. Surveillance and Security
The advancements in vision-based AI and IoT-connected cameras have heightened the potential for AI in cybersecurity (Pazho et al., 2023), including in Cyber Security Incident Response Teams (CSIRTs) (Killcrece et al., 2003). Determining the optimal level of AI autonomy in these settings is critical. A study by Hauptman et al. (Hauptman et al., 2023) explored this through a mixed-method approach, assessing AI’s role across the CSIRT cycle—preparation, identification, containment, eradication, and recovery. Surveying 103 cybersecurity professionals and conducting interviews with 22, the study found that higher AI autonomy is preferred during the identification phase, while lower autonomy suits high-risk phases better. This suggests that adaptive AI, with proper team training, can integrate effectively into CSIRTs, enhancing teamwork and response efficiency. Furthermore, the real-time update preference and successful integration of human and machine intelligence in crowd-powered camera sensing, as noted by Guo et al. (Guo et al., 2018), highlights the potential for enhancing cybersecurity measures through improved situational awareness and resource management.
Moreover, the utilization of LLMs for video surveillance analysis, as proposed by Chen et al. (Chen et al., 2023), with the VideoLLM framework, demonstrates the potential for transferring LLMs’ understanding and reasoning capabilities to video understanding tasks, offering a unified approach for various video-based security challenges. Additionally, Jain’s (Jain, 2023) insights into integrating ChatGPT with video security underscore the gradually increasing impact of foundation models on physical security systems, enhancing operational efficiency and surfacing new insights for businesses.
5.4. Games
Games serve as a dynamic platform for interaction, typically involving multiple human players engaging through various forms of competition or collaboration, each game offering unique goals and gameplay mechanics. The essence of these interactions often relies on the expressiveness unique to human participants, critical for optimal play. However, the advent of advanced Large Language Models (LLMs), such as ChatGPT, has expanded the potential for AI to mimic this human expressiveness, opening new avenues for AI to compete alongside or against humans in a broader array of gaming contexts. This segment delves into the integration of LLM-driven AIs as alternatives to human players within games, alongside their application in generating content for video games, marking a significant evolution in game design and interaction.
5.4.1. LLM-Based AI as Players
Artificial intelligence has long been utilized in video games, from serving as basic adversaries in titles like Mario Kart (Switch, 2017) to mastering complex games such as chess, demonstrating AI’s algorithmic prowess. Yet, these AIs often fell short as complete substitutes for human players due to their lack of flexibility and limited language processing capabilities. The advent of LLMs has shifted this landscape, offering AIs that can engage with humans in more expressive ways, suitable across a broader spectrum of games.
Research into LLMs as game players, while in its nascent stages, has yielded intriguing insights. For instance, Frans (Frans, 2021) showcases "AI Charades" using GPT-2 to illustrate LLMs’ potential in games requiring deep language interaction, highlighting their structured uncertainty—an attribute highly valued in games traditionally reliant on human partners. Xu et al.’s (Xu et al., 2023) exploration into LLMs within the social deduction game Werewolf demonstrates these AIs’ capability for strategic play, underscoring the necessity for advanced language understanding and social intuition. However, challenges such as limited input lengths and the impracticality of fine-tuning LLMs for specific games highlight current limitations.
Further critical examination by Sobieszek and Price (Sobieszek and Price, 2022) raise concerns over LLMs’ reliability in information delivery, emphasizing the risk of information pollution due to GPT-3’s occasionally untruthful responses. Research evaluating GPT models in social and moral contexts reveals mixed results, with models showing promise in single-player strategies but struggling in cooperative, ethical scenarios (Akata et al., 2023). This suggests a need for further development in LLMs’ ability to navigate complex social interactions and ethical decision-making. Despite these challenges, the expressive capabilities of LLMs compared to earlier AI models mark a significant advancement in human-computer interaction within games.
5.4.2. LLMs for Content Generation in Games
Large Language Models have significantly impacted game development, from generating complex game levels as demonstrated by Todd et al.(Todd et al., 2023), to enriching RPGs with dynamic quest narratives through advancements shown by Vartinen et al. (Värtinen et al., 2022). These models enhance game narratives and environment design, also facilitating interactive tutorials and support systems, a concept explored by Frans (Frans, 2021) with Language Model Games. Despite their potential, challenges persist, including ensuring content quality and coherence, addressing ethical concerns, and maintaining player immersion. The instructRL framework by Hu and Sadigh (Hu and Sadigh, 2023) offer insights into fine-tuning LLM integration for effective human-AI interaction in games. Advancing LLM applications in gaming requires addressing these hurdles through hybrid approaches and ethical guidelines.
5.5. Education
The integration of AI in education is revolutionizing teaching and learning by enabling personalized learning experiences through the analysis of student performance and real-time feedback. This advancement fosters increased student engagement and improved learning outcomes, yet raises concerns about data privacy and the transparency of AI decisions.
Intelligent Tutoring Systems (ITSs), such as Carnegie Learning’s Mika platform, demonstrate AI’s ability to tailor education to individual needs, showing promising results like higher exam scores and reduced dropout rates (Nwana, 1990; VanLehn, 2011). However, the opacity of AI’s decision-making has sparked debates over potential biases (Whitaker et al., 2013). Despite these challenges, companies like Century Tech and Fishtree have successfully incorporated AI to enhance personalized learning experiences (Kasinathan et al., 2017). Virtual learning agents, including Georgia Tech’s chatbot Jill Watson, have been developed to supplement teaching by handling administrative queries or facilitating peer collaboration, albeit with risks of fostering parasocial relationships with digital avatars (Eicher et al., 2018; Hartle and Kaczorowski, 2019). Moreover, automated essay scoring systems offered by EdX and Pearson exemplify AI’s efficiency in grading, aligning closely with human grader assessments while saving time and costs (Dikli, 2006). Nonetheless, there remains skepticism about AI’s ability to fully grasp the subtleties of written expressions, highlighting the ongoing need for advancements in natural language processing.
The use of LLMs in education is poised to personalize learning experiences further. Extance (Extance, 2023) explores the practical deployment of AI chatbots as interactive learning tools, demonstrating their potential to enhance the teaching and learning dynamic significantly. However, Milano et al. (Silvia Milano, 2023) raise critical points regarding the operational, ethical, and environmental considerations of LLMs in educational settings, emphasizing the need for sustainable and transparent practices. These insights offer a deeper look at AI’s role in education, balancing technological innovation with mindful consideration of its broader impacts (Wang and Demszky, 2023; Hellas et al., 2023; Chang et al., 2023).
5.6. Accessibility
Assistive technologies powered by artificial intelligence (AI) are revolutionizing support for individuals with sensory and physical disabilities, offering innovative solutions for improved mobility and communication. A notable advancement is a deep learning-based navigation assistant for the visually impaired, which employs computer vision to facilitate obstacle detection and navigation with remarkable accuracy, as demonstrated in the work by (Kumar and Jain, 2022). Similarly, the development of communication aids for the deaf, utilizing audio-visual recognition algorithms, provides an automated solution that enhances daily interactions, a concept explored by (Ozarkar et al., 2020). These technologies exemplify the transformative potential of AI in creating more inclusive environments, with significant improvements in user independence and quality of life. Further contributions include a compact, low-cost visual aid system for blind individuals that integrates advanced image processing for obstacle avoidance and object detection (Khan et al., 2020), and an AI-enabled sign language recognition system that supports sentence-level communication for the hearing/speech impaired (Wen et al., 2021). Additionally, the deployment of service robots to assist the elderly in shopping tasks demonstrates the practical applications of AI in facilitating daily activities (Ghazal et al., 2021).
In addition to these technological innovations, recent studies highlight the opportunities and challenges presented by large language models in promoting accessibility. Taheri et al.(Taheri et al., 2023) explore the potential of text-to-image generation models to revolutionize accessible art creation, demonstrating how generative text models can reduce typing effort and make visual art creation more accessible to people with diverse abilities. In contrast, Gadiraju et al.(Gadiraju et al., 2023) shed light on the subtle yet harmful biases LLMs can perpetuate towards disabled communities, emphasizing the importance of incorporating disability-positive resources in training data to foster a more inclusive digital environment. These insights underline the evolving landscape of AI-driven accessibility tools, while also stressing the need for continuous refinement and inclusivity in their development.
6. Conclusion
This survey on Human-AI (HAI) Teaming, with a focus on Large Pre-trained Models (LPtM), underscores a notable evolution in collaborative problem-solving and decision-making. Integrating human expertise into AI model training has enhanced performance and fostered intuitive and effective human-AI interactions. Safety, security, and trust emerge as critical components, requiring a balanced approach between technological progress and ensuring privacy and fairness. The survey’s exploration across various sectors, including healthcare, autonomous vehicles, and education, showcases HAI’s versatility in addressing unique challenges and enhancing societal benefits. As we advance, the integration of human intelligence with LPtMs looks like a promising domain in unlocking innovative solutions and improving overall decision-making. The future of HAI Teaming will hinge on navigating its challenges responsibly, ensuring ethical alignment, and maximizing its societal impact. This research contributes to the academic dialogue, offering insights for future exploration, policy-making, and strategic development towards a collaborative, effective, and equitable HAI ecosystem.